Let β„’{𝑓(𝑑)} = 𝐹(𝑠). Show that 𝑓(𝑑) = βˆ’ 1 𝑑 β„’ βˆ’1 {𝐹 β€² (𝑠)} Thus, if we know how to invert 𝐹 β€² (𝑠) then we know how to invert 𝐹(𝑠). Use this information to find the Laplace inverse transform of (i) arctan ( π‘Ž 𝑠 ), (ii) ln ( 𝑠+π‘Ž π‘ βˆ’π‘Ž ).Β 


1
Expert's answer
2021-10-01T04:34:27-0400

"F(s)=\int^{\infin}_0e^{-ts}f(t)dt"


"F'(s)=e^{-ts}f(t)"

"L^{-1}(F(s))=f(t)"

"L^{-1}(F'(s))=f(t)\/(-t)"

"f(t)=-tL^{-1}(F'(s))"


i)

"F'(s)=\frac{a}{1+s^2}"

"f(t)=-atsint"


ii)

"F'(s)=\frac{1}{s+a}-\frac{1}{s-a}"

"f(t)=-t(e^{-at}-e^{at})"